OpenStudy (anonymous):

Evaluate the following expression. 6^-3

OpenStudy (mle):

the negative means that you flip it over for example 3^ -1 = 1/3 so 6^-3 = 1 / (6^3) then we should find 6^3 that means 6 * 6 * 6

OpenStudy (anonymous):

Oh okay!

OpenStudy (anonymous):

Please leave me alone..

OpenStudy (mle):

If you get suck, just say so and I'll break it down for you more :) If you get the answer, tell me and I'll help check it for you ^_^

OpenStudy (anonymous):

Thank you!!!!